# Helping or Herding? Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

### v1

@@ Abstract
Reward models play a key role in aligning language model applications towards human preferences. However, this setup creates an incentive for the language model to exploit errors in the reward model to achieve high estimated reward, a phenomenon often termed reward hacking. A natural mitigation is to train an ensemble of reward models, aggregating over model outputs to obtain a more robust reward estimate. We explore the application of reward ensembles to alignment at both training time (through reinforcement learning) and inference time (through reranking). First, we show that reward models are underspecified: reward models that perform similarly in-distribution can yield very different rewards when used in alignment, due to distribution shift. Second, underspecification can result in overoptimization, where alignment to one reward model does not improve reward as measured by another reward model trained on the same data. Third, overoptimization is mitigated by the use of reward ensembles, and ensembles that vary by their pretraining seeds generalize better than ensembles that differ only by their fine-tuning seeds. However, even pretrain reward ensembles do not eliminate reward hacking: we show several qualitative reward hacking phenomena that are not mitigated by ensembling because all reward models in the ensemble exhibit similar error patterns.
@@ 1 Introduction
Motivated by these findings, we systematically investigate reward model ensembles as a possible remedy for reward hacking. Assuming different models err in different ways, ensembling can leverage reward uncertainty across the ensemble during alignment (see Figure˜1, Left). We explore several techniques for aggregating scores across the ensemble, e.g., taking the median score as a robust estimate of the true reward of the policy. We also consider two types of ensembles: pretrain ensembles, where different members of the ensemble differ in the random seed used during the pretraining phase, and finetune ensembles, where members differ only in the random seed used during finetuning. These ensembles are then evaluated across several types of policies and preference annotations: dialogue preferences for a helpful assistant (Bai et al., 2022), summarization quality (Stiennon et al., 2020), and whether a single-document summary is grounded in its source text (Roit et al., 2023).
We find that pretrain ensembles substantially outperform finetune ensembles. Moreover, they consistently outperform single reward models, unlike finetune ensembles, which in many cases are comparable to single reward models. However, our analysis also reveals that policies trained with ensembles are still susceptible to reward hacking: different reward models sometimes share similar error patterns, which in turn propagate to the ensemble (see Figure˜1, Right). This is exploited and amplified by the policy, leading, for example, to outputs that are too short when tuning for factuality, too verbose when tuning for summarization quality, or responses that follow a particular format that is often unsuitable, when training a helpful assistant. Thus, it is possible that methods that, unlike ensembles, are aware of the distance of outputs from the reward data distribution (Liu et al., 2020) could provide more reliable estimates of uncertainty.
@@ 2.3 Experimental Setup
TL;DR I’ll try to keep this short! **Background** * I’ve always been an on again/off again (very casual!) jogger, typically doing 3 - 5 k * My knees have always been finicky, and I went to a physio who thought I had “runner’s knee” […] I had my baby a year ago, so all in all I haven’t run for about 1.5 years […] However, when I run even the tinyiest bit, or even go on a long walk or a hike, my pelvis gets very sore and tight […] I think pregnancy messed with my body, now I can’t even run even the smallest amount without pain in my pelvis and knees. I’m fairly certain the problem isn’t just that I’m completely out of shape.
Helpfulness Human: Assistant, can you help me find books that have really popular antagonists? Assistant: Sure! I know many books that feature good antagonists, such as those in the Lord of the Rings books. Human: Who specifically from Lord of the Rings? Assistant: Well, Sauron, for example. If you want, I can explain who he was.
XSum/NLI The ex-Reading defender denied fraudulent trading charges relating to the Sodje Sports Foundation - a charity to raise money for Nigerian sport. Mr Sodje, 37, is jointly charged with elder brothers Efe, 44, Bright, 50 and Stephen, 42. Appearing at the Old Bailey earlier, all four denied the offence. The charge relates to offences which allegedly took place between 2008 and 2014. Sam, from Kent, Efe and Bright, of Greater Manchester, and Stephen, from Bexley, are due to stand trial in July. They were all released on bail. Former Premier League footballer Sam Sodje has appeared in court alongside three brothers accused of charity fraud.
@@
We will examine the performance of reward models (both single models and ensembles) across three tasks. An example from each task is provided in Table˜1.
@@
In BoN reranking, we rerank sampled sets of size $n\in\{2^{1},2^{2},\ldots,2^{5}\}$ for helpfulness and $\{2^{1},\ldots,2^{6}\}$ for tl;dr. Larger sets lead to higher reward at a cost of more expensive inference and larger deviation from $\pi_{\text{sft}}$. In RLHF, we obtain a trade-off between the KL from $\pi_{\text{sft}}$ and the expected reward by training multiple times, varying the value of $\lambda$. Low values of $\lambda$ correspond to high KL and high reward, while high values of $\lambda$ entail low KL and low reward. For each value of $\lambda$ we train roughly to convergence using a predetermined fixed number of steps (all hyperparameter values, including $\lambda$ and the number of steps, are in Appendix˜C). Coste et al. (2023) trade-off KL and reward by tracking their values during training; however, for any particular value of KL the reward might still be underoptimized during training (i.e., there can exist a different policy $\pi(y\mid x)$ with better reward, but the same $\mathrm{KL}(\pi(y\mid x)\|\pi_{\mathrm{sft}}(y\mid x))$, which can be found with longer training).
@@
We use two metrics to quantify generalization of reward models—reward by a larger model and win rate. Similar to past work (Gao et al., 2023; Coste et al., 2023), we use a larger reward model to evaluate the generalization of models trained with a smaller reward model. We train a T5-XXL reward model by taking the publicly available T5-XXL (Raffel et al., 2020) and finetuning it as described above. Table˜2 details the performance of reward models of different sizes on the three tasks, and it can be seen that T5-XXL outperforms the best T5-XL model. We report both average reward of the T5-XXL evaluator as well as win rate, which is the fraction of prompts for which the response sampled from the aligned policy $\pi$ has higher reward compared to $\pi_{\text{sft}}$.
The errors of the T5-XXL autoeval model might correlate with errors of the smaller T5 models because they are trained on the same preference data. For this reason, we also evaluate win rate according to a prompted PALM-2 model which was not exposed to the reward training data but was instruction-tuned on FLAN (Wei et al., 2022). Given a prompt $x$, we sample a response $y_{\text{sft}}$ from $\pi_{\text{sft}}$ and $y_{\text{rlhf}}$ from $\pi$. We then ask PALM-2 which response is better, using a hand-engineered prompt proposed by Rafailov et al. (2023). To avoid position bias we run PALM-2 on the two possible orderings $(y_{\text{sft}},y_{\text{rlhf}})$ and $(y_{\text{sft}},y_{\text{rlhf}})$, sample $K=8$ outputs for each order and determine the winner on this prompt through majority voting. This style of evaluation has become common recently (Dubois et al., 2023; Singhal et al., 2023) and was shown to correlate well with human judgements (Rafailov et al., 2023).
@@ 3 Underspecification in Reward Models
We now analyze alignment strategies that use a single reward model, and demonstrate that reward models are underspecified. First, Table˜2 shows the average in-distribution accuracy across the 25 different reward models, together with the standard deviation (which is low in-distribution).
@@
The story changes, however, when we move to out-of-distribution data. Figure˜2 shows the expected reward achieved by BoN as a function of the number of sampled candidates, $n$, for three reward model scales (KL is approximately $\log n-\frac{n-1}{n}$). The dotted green line shows the expected reward of the top-ranked output according to the reranker itself, while the dashed orange line shows the expected reward of the same output according to reward models that share a pretrain seed. The solid blue line shows the expected reward according to reward models that do not share a pretrain seed. Unsurprisingly, the reranker scores its own top outputs more favorably than the other reward models do. However, the reranker’s outputs are scored significantly less favorably by reward models which do not share a pretrain with the ranker. Reward models that share a pretrain seed with the ranker model overestimate the true reward of the top-ranked output—suggesting that finetune ensembles are not sufficiently diverse because of the shared pretraining state of each of the ensemble’s members. Notably, this gap does not disappear with scale, and is present for base, large, and XL models.
@@
Moving to alignment, differences in estimated rewards induce different policies from the BoN strategy: Figure˜3 shows the effects on agreement of the top-ranked summary when reward models do (crosses) or do not (circles) share pretraining seeds. Different reward models tend to produce different 1-best outputs. Again these differences are strongly associated with the pretraining seed: for example, two reward models from different pretrains will choose a different best-of-16 output more than half the time for both tl;dr and helpfulness and in all scales.
@@
Last, Figure˜4 analyzes the evolution of agreement of the estimated reward scores when performing RLHF on tl;dr for reward models of various scales. Specifically, we align a policy using a single reward model, and then measure how well pairs of reward models agree on the ranking of samples from that policy using Spearman rank correlation. To compute Spearman, we sample 5 completions for each prompt in the validation set from a policy model, at 2K step intervals during RLHF. We compare the agreement between a set of 5 reward models that share the same pre-training seed and a set of 5 that do not (both sets include the reward model used to drive RLHF). For each prompt, we compute Spearman correlation across all ten pairs in each set and report the mean correlation over the pairs. The correlation of models that do not share a pretrain is lower compared to models that share a pretrain seed. Moreover, correlation goes down during RLHF, indicating that the uncertainty about the true reward increases as a result of alignment.
@@ 4.2 Experiments
We now evaluate reward model ensembles across all tasks. Figure˜5 shows the results of ensembling in best-of-$n$ reranking, as measured by an XXL-scale fine-tuned reward model. Pretrain ensembles consistently improve performance over individual reward models, especially for higher values of $n$ for both tl;dr and helpfulness. Finetune ensembles, conversely, improve performance in some cases and are comparable in others. For example, on tl;dr a pretrain ensemble with the mean aggregator achieves a win rate of 90% over the SFT outputs at the XL scale, while the win rate of a finetune ensemble with the same mean aggregator is 87.3%. The win rate of the average individual XL-scale reward model is 85.3% (see Table˜7). For visual clarity, in Figure˜5 we show only two aggregators: mean and mean_minus_std; see Appendix˜A for results with other aggregators. In general, the differences between aggregators are small, with mean usually performing at, or near, the top. More conservative aggregators (min and mean_minus_std) come out slightly ahead of mean at the smaller scales on tl;dr, suggesting that high variance may be a bigger issue in this setting.
Figure˜6 shows the KL-reward trade-off of ensemble reward models in RLHF for tl;dr and helpfulness (evaluated with the finetuned T5-XXL model). In such plots, a better model is one that improves reward and/or reduces the value of KL from the original SFT policy (Gao et al., 2023; Coste et al., 2023). Indeed, similar to BoN, pretrain ensembles consistently outperform both finetune ensembles as well as the average individual model. We present results for the median and mean aggregators for visual clarity, and report full numerical results in Appendix˜B. In RLHF, KL values are much higher than BoN (which is bounded by $\approx 3.17$ for $n=64$). Consequently, in this setting we witness explicit reward hacking, in which the T5-XXL rewards decrease even as the RLHF objective improves. This happens most prominently for individual models, in many cases for finetune ensembles, and most rarely for pretrain ensembles—where T5-XXL reward scores decrease only when RLHF uses a T5-Base reward model. Thus, our experiments on real data yield more negative conclusions than Coste et al. (2023) about the potential of ensembles to eliminate reward overoptimization.
@@
Because the T5-XXL autoeval model is trained on the same data distribution as the reward models used for best-of-$n$ and RLHF, it may overstate their performance. For this reason, we also use a zero-shot autoeval model (Palm-2), as described in Section˜2.3.222To reduce costs, we used Palm-2-M, except for RLHF on helpfulness, where we evaluated with Palm-2-L. In best-of-$n$, we apply this model to $n=64$ for tl;dr and $n=32$ for helpfulness, focusing on the XL-scale reward models. The results are generally in accord with Figure˜5. On tl;dr, the pretrain ensemble win rate is 82.3%; finetune ensemble: 79.7%; average single reward: 77.5%. These differences are all statistically significant. On helpfulness, the pretrain ensemble win rate is 81.3%; finetune ensemble, 81.8%; average single ensemble, 77.5%. In this case, the difference between the two ensembles is not significant ($p\approx.3$), but the difference between each ensemble and the single reward model is significant at $p<.01$. These win rates are similar to those of the T5-XXL autoeval model (78-82%, see Table˜8).
Figure 7: Pretrain ensemble reward models lead to more favorable win rate vs. KL tradeoffs, as judged by a prompted FLAN Palm-2 for helpfulness (Left) and tl;dr (Right) in RLHF with T5-XL reward model.
Figure˜7 shows win rates in RLHF alignment for helpfulness (Left) and tl;dr (Right) with Palm-2. Overall, these results mirror those of the T5-XXL evaluation—pretrain ensembles outperform finetune ensembles, especially in helpfulness. Interestingly, reward hacking is evident for lower values of KL in helpfulness, which suggests that evaluating with a model trained on the preference data overestimates performance.
@@
Figure˜8 shows RLHF results for xsum/nli. Here we see a relatively small improvement for ensembles compared to individual models, and a very small difference between pretrain and finetune ensembles. We conjecture this is because xsum/nli optimizes for a particular aspect of the response, namely its factuality. This allows all models to find simple and similar strategies that lead to high reward (for example, emitting short responses with limited content), and thus ensembling does not lead to large gains in performance. We further elaborate on this when discussing limitations of ensembles in §5.

### v2

# Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

@@ Abstract
Reward models play a key role in aligning language model applications towards human preferences. However, this setup creates an incentive for the language model to exploit errors in the reward model to achieve high estimated reward, a phenomenon often termed reward hacking. A natural mitigation is to train an ensemble of reward models, aggregating over model outputs to obtain a more robust reward estimate. We explore the application of reward ensembles to alignment at both training time (through reinforcement learning) and inference time (through reranking). First, we show that reward models are underspecified: reward models that perform similarly in-distribution can yield very different rewards when used in alignment, due to distribution shift. Second, underspecification results in overoptimization, where alignment to one reward model does not improve reward as measured by another reward model trained on the same data. Third, overoptimization is mitigated by the use of reward ensembles, and ensembles that vary by their pretraining seeds lead to better generalization than ensembles that differ only by their fine-tuning seeds, with both outperforming individual reward models. However, even pretrain reward ensembles do not eliminate reward hacking: we show several qualitative reward hacking phenomena that are not mitigated by ensembling because all reward models in the ensemble exhibit similar error patterns.
@@ 1 Introduction
Motivated by these findings, we systematically investigate reward model ensembles as a possible remedy for reward hacking. Assuming different models err in different ways, ensembling can leverage reward uncertainty across the ensemble during alignment (see Figure 1, Left). We explore several techniques for aggregating scores across the ensemble, e.g., taking the median score as a robust estimate of the true reward of the policy. We also consider two types of ensembles: pretrain ensembles, where different members of the ensemble differ in the random seed used during the pretraining phase, and finetune ensembles, where members differ only in the random seed used during finetuning. These ensembles are then evaluated across several types of policies and preference annotations: dialogue preferences for a helpful assistant (Bai et al., 2022), summarization quality (Stiennon et al., 2020), and whether a single-document summary is grounded in its source text (Roit et al., 2023).
We find that pretrain ensembles substantially outperform finetune ensembles. Moreover, they consistently outperform single reward models, unlike finetune ensembles, which in many cases are comparable to single reward models. However, our analysis also reveals that policies trained with ensembles are still susceptible to reward hacking: different reward models sometimes share similar error patterns, which in turn propagate to the ensemble (see Figure 1, Right). This is exploited and amplified by the policy, leading, for example, to outputs that are too short when tuning for factuality, too verbose when tuning for summarization quality, or responses that follow a particular format that is often unsuitable, when training a helpful assistant. Thus, it is possible that methods that, unlike ensembles, are aware of the distance of outputs from the reward data distribution (Liu et al., 2020) could provide more reliable estimates of uncertainty.
@@ 2.3 Experimental Setup
TL;DR
I’ll try to keep this short! **Background** * I’ve always been an on again/off again (very casual!) jogger, typically doing 3 - 5 k * My knees have always been finicky, and I went to a physio who thought I had “runner’s knee” […] I had my baby a year ago, so all in all I haven’t run for about 1.5 years […] However, when I run even the tinyiest bit, or even go on a long walk or a hike, my pelvis gets very sore and tight […]
I think pregnancy messed with my body, now I can’t even run even the smallest amount without pain in my pelvis and knees. I’m fairly certain the problem isn’t just that I’m completely out of shape.
Helpfulness
Human: Assistant, can you help me find books that have really popular antagonists? Assistant: Sure! I know many books that feature good antagonists, such as those in the Lord of the Rings books. Human: Who specifically from Lord of the Rings? Assistant:
Well, Sauron, for example. If you want, I can explain who he was.
XSum/NLI
The ex-Reading defender denied fraudulent trading charges relating to the Sodje Sports Foundation - a charity to raise money for Nigerian sport. Mr Sodje, 37, is jointly charged with elder brothers Efe, 44, Bright, 50 and Stephen, 42. Appearing at the Old Bailey earlier, all four denied the offence. The charge relates to offences which allegedly took place between 2008 and 2014. Sam, from Kent, Efe and Bright, of Greater Manchester, and Stephen, from Bexley, are due to stand trial in July. They were all released on bail.
Former Premier League footballer Sam Sodje has appeared in court alongside three brothers accused of charity fraud.
@@
We will examine the performance of reward models (both single models and ensembles) across three tasks. An example from each task is provided in Table 1.
  • [leftmargin=*]
@@
In BoN reranking, we rerank sampled sets of size $n\in\{2^{1},2^{2},\ldots,2^{5}\}$ for helpfulness and $\{2^{1},\ldots,2^{6}\}$ for tl;dr. Larger sets lead to higher reward at a cost of more expensive inference and larger deviation from $\pi_{\text{sft}}$. In RLHF, we obtain a trade-off between the KL from $\pi_{\text{sft}}$ and the expected reward by training multiple times, varying the value of $\lambda$. Low values of $\lambda$ correspond to high KL and high reward, while high values of $\lambda$ entail low KL and low reward. For each value of $\lambda$ we train roughly to convergence using a predetermined fixed number of steps (all hyperparameter values, including $\lambda$ and the number of steps, are in Appendix C). Coste et al. (2023) trade-off KL and reward by tracking their values during training; however, for any particular value of KL the reward might still be underoptimized during training (i.e., there can exist a different policy $\pi(y\mid x)$ with better reward, but the same $\mathrm{KL}(\pi(y\mid x)\|\pi_{\mathrm{sft}}(y\mid x))$, which can be found with longer training).
@@
We use two metrics to quantify generalization of reward models—reward by a larger model and win rate. Similar to past work (Gao et al., 2023; Coste et al., 2023), we use a larger reward model to evaluate the generalization of models trained with a smaller reward model. We train a T5-XXL reward model by taking the publicly available T5-XXL (Raffel et al., 2020) and finetuning it as described above. Table 2 details the performance of reward models of different sizes on the three tasks, and it can be seen that T5-XXL outperforms the best T5-XL model. We report both average reward of the T5-XXL evaluator as well as win rate, which is the fraction of prompts for which the response sampled from the aligned policy $\pi$ has higher reward compared to $\pi_{\text{sft}}$.
The errors of the T5-XXL autoeval model might correlate with errors of the smaller T5 models because they are trained on the same preference data. For this reason, we also evaluate win rate according to a prompted PaLM-2-Large model, which was not exposed to the reward training data but was instruction-tuned on FLAN (Wei et al., 2022). Given a prompt $x$, we sample a response $y_{\text{sft}}$ from $\pi_{\text{sft}}$ and $y_{\text{rlhf}}$ from $\pi$. We then ask PaLM-2 which response is better, using a hand-engineered prompt proposed by Rafailov et al. (2023). To avoid position bias we run PaLM-2 on the two possible orderings $(y_{\text{sft}},y_{\text{rlhf}})$ and $(y_{\text{sft}},y_{\text{rlhf}})$, sample $K=8$ outputs for each order and determine the winner on this prompt through majority voting. This style of evaluation has become common recently (Dubois et al., 2023; Singhal et al., 2023) and was shown to correlate well with human judgements (Rafailov et al., 2023).
@@ 3 Underspecification in Reward Models
We now analyze alignment strategies that use a single reward model, and demonstrate that reward models are underspecified. First, Table 2 shows the average in-distribution accuracy across the 25 different reward models, together with the standard deviation (which is low in-distribution).
@@
The story changes, however, when we move to out-of-distribution data. Figure 2 shows the expected reward achieved by BoN as a function of the number of sampled candidates, $n$, for three reward model scales (KL is approximately $\log n-\frac{n-1}{n}$). The dotted green line shows the expected reward of the top-ranked output according to the reranker itself, while the dashed orange line shows the expected reward of the same output according to reward models that share a pretrain seed. The solid blue line shows the expected reward according to reward models that do not share a pretrain seed. Unsurprisingly, the reranker scores its own top outputs more favorably than the other reward models do. However, the reranker’s outputs are scored significantly less favorably by reward models which do not share a pretrain with the ranker. Reward models that share a pretrain seed with the ranker model overestimate the true reward of the top-ranked output—suggesting that finetune ensembles are not sufficiently diverse because of the shared pretraining state of each of the ensemble’s members. Notably, this gap does not disappear with scale, and is present for base, large, and XL models.
@@
Moving to alignment, differences in estimated rewards induce different policies from the BoN strategy: Figure 3 shows the effects on agreement of the top-ranked summary when reward models do (crosses) or do not (circles) share pretraining seeds. Different reward models tend to produce different 1-best outputs. Again these differences are strongly associated with the pretraining seed: for example, two reward models from different pretrains will choose a different best-of-16 output more than half the time for both tl;dr and helpfulness and in all scales.
@@
Last, Figure 4 analyzes the evolution of agreement of the estimated reward scores when performing RLHF on tl;dr for reward models of various scales. Specifically, we align a policy using a single reward model, and then measure how well pairs of reward models agree on the ranking of samples from that policy using Spearman rank correlation. To compute Spearman, we sample 5 completions for each prompt in the validation set from a policy model, at 2K step intervals during RLHF. We compare the agreement between a set of 5 reward models that share the same pre-training seed and a set of 5 that do not (both sets include the reward model used to drive RLHF). For each prompt, we compute Spearman correlation across all ten pairs in each set and report the mean correlation over the pairs. The correlation of models that do not share a pretrain is lower compared to models that share a pretrain seed. Moreover, correlation goes down during RLHF, indicating that the uncertainty about the true reward increases as a result of alignment.
@@ 4.2 Experiments
We now evaluate reward model ensembles across all tasks. Figure 5 shows the results of ensembling in best-of-$n$ reranking, as measured by an XXL-scale fine-tuned reward model. Pretrain ensembles consistently improve performance over individual reward models, especially for higher values of $n$ for both tl;dr and helpfulness. Finetune ensembles, conversely, improve performance in some cases and are comparable in others. For example, on tl;dr a pretrain ensemble with the mean aggregator achieves a win rate of 90% over the SFT outputs at the XL scale, while the win rate of a finetune ensemble with the same mean aggregator is 87.3%. The win rate of the average individual XL-scale reward model is 85.3% (see Table 7). For visual clarity, in Figure 5 we show only two aggregators: mean and mean_minus_std; see Appendix A for results with other aggregators. In general, the differences between aggregators are small, with mean usually performing at, or near, the top. More conservative aggregators (min and mean_minus_std) come out slightly ahead of mean at the smaller scales on tl;dr, suggesting that high variance may be a bigger issue in this setting.
Figure 6 shows the KL-reward trade-off of ensemble reward models in RLHF for tl;dr and helpfulness (evaluated with the finetuned T5-XXL model). In such plots, a better model is one that improves reward and/or reduces the value of KL from the original SFT policy (Gao et al., 2023; Coste et al., 2023). Indeed, similar to BoN, pretrain ensembles consistently outperform both finetune ensembles as well as the average individual model. We present results for the median and mean aggregators for visual clarity, and report full numerical results in Appendix B. In RLHF, KL values are much higher than BoN (which is bounded by $\approx 3.17$ for $n=64$). Consequently, in this setting we witness explicit reward hacking, in which the T5-XXL rewards decrease even as the RLHF objective improves. This happens most prominently for individual models, in many cases for finetune ensembles, and most rarely for pretrain ensembles—where T5-XXL reward scores decrease only when RLHF uses a T5-Base reward model. Thus, our experiments on real data yield more negative conclusions than Coste et al. (2023) about the potential of ensembles to eliminate reward overoptimization.
@@
[…]
(a) BoN + tl;dr
(b) BoN + helpfulness
(c) RLHF + tl;dr
(d) RLHF + helpfulness
Figure 7: Using a prompted autoevaluator (PaLM-2-FLAN), ensemble reward models offer significantly better win rates on both tl;dr and helpfulness. Here all reward models are XL-scale.
@@
Figure 8 shows RLHF results for xsum/nli. Here we see a relatively small improvement for ensembles compared to individual models, and a very small difference between pretrain and finetune ensembles. We conjecture this is because xsum/nli optimizes for a particular aspect of the response, namely its factuality. This allows all models to find simple and similar strategies that lead to high reward (for example, emitting short responses with limited content), and thus ensembling does not lead to large gains in performance. We further elaborate on this when discussing limitations of ensembles in §5.